# Reasoning Grasping via Multimodal Large Language Model

### v1

@@ I Introduction
Robotic grasping has long been a subject of extensive study. The utilization of CNN-based neural networks has demonstrated efficiency in generating high-quality grasping poses from visual input [32, 22, 19, 24, 34, 28]. One significant limitation of these methods is the lack of scene understanding, resulting in robots being unable to identify the objects they are grasping. Recent developments have introduced methods that interpret language instructions such as “I want a stapler” [7] or “Pick the food box in front of the ball” [43] for targeted grasping, enabling robots to pick objects as directed. However, those methods will struggle when dealing with implicit instructions, where the desired object is not explicitly named. For example, if a user says “I need to drink water”, an intelligent assistant robot should autonomously identify and grasp a cup from a selection of household items on the table, although “cup” does not appear in the user’s instruction. Such scenarios, prevalent in real-world setting, underscore the importance of advanced understanding and autonomy in robotic grasping — an area that remains a significant challenge.
@@
Recently, the advance of large language models like ChatGPT [33] has introduced new possibilities for language reasoning, significantly enhancing our ability to process and interpret complex language structures. Beyond text, numerous studies have leveraged multi-modal Large Language Models (LLMs) for enhanced visual understanding and interaction [51, 12, 25], enabling these models to interpret and generate responses based on a combination of textual and visual inputs. However, these approaches have been confined largely to visual comprehension and text generation, lacking the capability to generate precise robotic actions. While there have been efforts to integrate LLMs with robotics [2, 17, 21, 37, 44, 18], these have predominantly focused on high-level planning or have modularized the LLM for reasoning while relying on traditional methods for downstream actions. Brohan et al. [4] utilize a Vision Language Model (VLM) to transfer web knowledge to robot control, but it requires extensive efforts to train on a large amount of data.
In this work, we introduce the novel task of Reasoning Grasping, where the robot is tasked with generating appropriate grasping poses based on implicit instructions. This task presents two significant challenges, primarily due to the nuanced requirements of interpreting complex, context-rich instructions and accurately generating the corresponding grasp poses.
@@
To address these challenges, we propose a model that is uniquely designed to directly output grasp poses according to visual scenes and complex instructions in cluttered environments. Specifically, this model initially leverages a multi-modal LLM to interpret both images and instructions, producing a textual output that specifies the grasping target. Given the LLM’s output tends to be verbose, here we employ a special token strategy to identify those tokens that are highly relevant to the grasping target (i.e., names of target objects). Subsequently, embeddings of these identified tokens are extracted and passed through a projection network to derive a textual feature. This feature, encapsulated with essential textual information about the target, is then incorporated into the grasping prediction framework to enhance the model’s ability to accurately predict grasp poses.
However, we observe that existing grasping datasets primarily concentrate on either only visual reasoning or object grasping based on explicit instructions. To address this deficiency, we introduce a comprehensive dataset specifically designed to support the development and benchmarking of models that tackle the challenge of reasoning grasping. This dataset encompasses a diverse array of visual scenes paired with complex, implicit instructions, meticulously tailored to train and assess the reasoning and grasping capabilities of robotic systems in cluttered environments.
@@
  • We introduce a novel task of reasoning grasping, which challenges the system to not only interpret complex and implicit instructions but also to accurately generate precise grasping poses for robotic execution. This is a crucial capability for truly intelligent robotic systems.
  • We propose a benchmark model for reasoning grasping that combines multi-modal LLMs with robotic grasping techniques. Our experiments demonstrate the model’s promising performance in generating precise grasp poses in response to language instructions.
  • We have developed a unique dataset for the reasoning grasping task. It includes 64 objects, 109 parts, 1730 reasoning instructions, and around 100 million grasping poses. This dataset provides a comprehensive resource for training and evaluating models on their ability to interpret and execute complex robotic grasping tasks based on various textual instructions.
@@ II-A Robotic Grasping
Robotic grasping has traditionally relied on analytical methods [29, 11, 36], which focus on understanding the geometry of objects or analyzing contact forces, and convolutional neural networks (CNNs) based methods [32, 22, 19, 24, 34, 28], which utilize large datasets of labeled grasping examples [20, 9, 14] to train models to predict grasps based on visual input. However, a critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions. They do not inherently know what they are grasping, which restricts their application in more dynamic, human-centric environments.
Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [27, 39, 41, 6, 47, 38, 46, 48, 50, 1]. This shift towards language-guided referring grasping allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios. For example, Tziafas et al. focus on referring grasp synthesis, predicting grasp poses for referred objects through natural language in cluttered scenes [43]. Chen et al. proposed to learn from visual and language features jointly and predict the 2D grasp box from the RGB images [7].
Despite advancements, a critical limitation persists in robotic datasets and methods: the dependence on explicit language commands for target identification. This restricts robots’ ability to understand nuanced, implicit instructions crucial for real-world interactions. Our dataset tackles this by incorporating reasoning tasks and detailed annotations for object parts and grasping poses, enabling robots to respond to complex instructions and more accurately align with human expectations.
@@ II-B LLMs for robotics
Recent significant breakthroughs in Large Language Models (LLMs) have been impressive [10, 5, 33, 42] in understanding and generating human-like text. The capabilities of LLMs extend well beyond text processing. Multi-modal LLMs can seamlessly integrate with other modalities, such as vision, enhancing their proficiency in tasks like visual understanding [25]. There are also a lot of efforts integrating LLMs with robotics. Many studies [2, 17, 21] have integrated LLMs into closed-loop planning structures, decomposing language-conditioned long-horizon tasks into small steps. Yet, the gap between language instructions and actions still remains. Furthermore, some studies [37, 44, 18] have employed program-like specifications to prompt LLMs, melding planning and action using a predefined library of action functions.
There is also an increase in research focused on utilizing LLMs specifically for robotic grasping. Tang et al. proposed to use LLMs to generate semantic knowledge and then selected task-oriented grasp pose from grasp candidates [40]. Mirjalili et al. employed Large Language Models to identify the optimal part of an object for grasping, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [30]. In contrast to the aforementioned methods, which utilize LLMs as just one component in a multi-module grasping framework and the grasp detection relies on another grasping module, we propose a more streamlined, end-to-end approach. Our model uniquely integrates vision-language capabilities, allowing it to process language instructions and directly output the appropriate grasping pose through conversational interaction. Moreover, our model can understand implicit instructions and is capable of engaging in multi-round conversations, enhancing its applicability and intelligence in robotic grasping tasks.
@@ III Reasoning Grasping
In this study, we introduce a novel problem of Reasoning Grasping, which involves predicting antipodal grasps for robots based on a combination of visual scenes and textual instructions. While the potential for integrating natural language into robotic manipulation is vast, its practical application in robotic grasping remains limited. Existing research [43, 46], utilizing techniques like the CLIP text encoder [35], has begun to address target-oriented grasping in cluttered environments based on straightforward language instructions, such as “Give me the banana.” However, these methods often fall short in real-world scenarios where human instructions are complex and implied, such as “I feel a little hungry.” This gap highlights the need for a more nuanced understanding and interpretation of natural language within robotic systems, driving the development of reasoning grasping.
The Reasoning Grasping task aims to generate a robotic grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$. The textural instruction $t$ is not limited to straightforward language instructions, it can also be complex and implicit instructions.
@@ IV Reasoning Grasping via Multi-modal LLMs
In this section, we present an innovative model for reasoning grasping, which combines conventional robotic grasping in [22] and reasoning capabilities from multi-modal LLMs. This method effectively bridges the gap between image-based grasp prediction techniques and the nuanced understanding inherent in textual language models. Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs to enhance and guide the robotic grasping prediction process.
This model addresses two key challenges: firstly, the adaptation of pre-trained multi-modal LLMs for efficient training in specific downstream tasks. While pre-trained LLMs possess strong reasoning abilities, their tendency to produce verbose and diverse outputs complicates their use in training other tasks, especially within an end-to-end framework. Secondly, the model emphasizes integrating these textual outputs from LLMs with image-based robotic grasping systems. This integration is crucial as it ensures that the rich linguistic context is effectively utilized to refine the subsequent grasping prediction.
@@ IV-A Architecture
The proposed model consists of two main components: the reasoning module and the grasping module, as illustrated in Fig. 1. In the reasoning module, we utilize the pre-trained LLaVA [25] enhanced with LoRA [16] fine-tuning. LLaVA integrates a CLIP [35] visual encoder with LLaMA [42], an open-source Large Language Model comparable in performance to GPT-3 [5]. And LoRA, known for its computational efficiency, is incorporated into both the projection layer and all linear layers of the LLMs.
@@
The key role of the LLMs in the reasoning module is to interpret instructions and accurately identify the target object or part for grasping. To counteract the typical verbosity of LLM outputs, we propose the use of a special token, i.e., [SPT]. This token is strategically placed around the names of the target object or the target part in the LLM’s textual outputs, i.e., [SPT] grasping target name [SPT]. The purpose of introducing this special token is to identify and isolate the crucial tokens, specifically, the names of the grasping target, from the broader textual outputs generated by LLMs.
@@
Once the grasping target is identified within the textual outputs using the special token, we proceed to extract the last-layer embeddings of its corresponding tokens. It is important to note that there could be multiple tokens associated with the identified grasping target, and we extract the embeddings for all corresponding tokens. These extracted embeddings encapsulate the essential linguistic information related to the grasping target. Subsequently, we compute the average of these embeddings denoted as $h_{avg}$, which is then passed through a projection network to obtain the feature $f$. This feature $f$ is then fed into the grasping module, guiding the prediction of the grasp pose.
@@
The grasping module, rooted in the CNN-based robotic grasping framework in [22], takes n-channel images as inputs. The input image is first passed through three convolutional layers and five residual layers. The extracted visual features are concatenated with the embeddings obtained from the reasoning module. Following this, we use the convolutional transpose operation to up-sample this concatenated feature, increasing its size to match the input image size. The model subsequently generates four output images, each representing different aspects of the grasp pose $g$: the grasp quality score $q$, the rotation angle expressed in $\cos 2\theta$ and $\sin 2\theta$, and the required width $w$ for the end-effector. The grasp central coordinate $(x,y)$ is determined from the output image of the grasp quality score. This is achieved by selecting the point in the image that corresponds to the maximum score.
The reasoning module processes inputs of RGB images $v$ and textual instructions $t$, while the grasping module utilizes RGB images $v$ and, optionally, depth images $v_{d}$, to accurately determine the grasp pose $g$. Specifically, the probability of generating the target grasp pose in the problem of reasoning grasping is formulated as
$p(g|v,v_{d},t)=p_{\theta_{r}}(f|v,t)\cdot p_{\theta_{g}}(g|v,v_{d},f),$ (1)
where $\theta_{r}$ and $\theta_{g}$ represent parameters in the reasoning module and the grasping module, respectively. And $f$ is the extracted feature derived from outputs of the multi-modal LLM, as illustrated in Fig. 1.
@@ IV-C Training Strategy
The proposed method leverages the pre-trained LLaVA model, with the LoRA technique for fine-tuning during the training phase. Notably, enhanced by LoRA find-tuning, the multi-modal LLM enables accurately producing textual outputs specifying the grasping target within just a few epochs. Based on this observation, our training strategy involves freezing the parameters of the LoRA-enhanced model after the initial epochs. Specifically, after the second epoch, we set the parameter $\lambda_{t}$ in the loss function to 00 and freeze the parameters of the LoRA-enhanced model. From the third epoch, our training focuses exclusively on other parameters outside the LoRA framework. This approach is designed to reduce the training time and expedite the convergence of the model, particularly in terms of grasping prediction.
@@
Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [14] with additional annotations for part-level grasping and accompanying reasoning instructions that specifically guide the grasping of corresponding objects.
@@ V Reasoning Grasping Dataset
Reasoning grasping represents a promising yet underexplored area in robotics, offering the potential to significantly enhance how robots interact with and understand their environments. Current research and datasets primarily focus on visual reasoning or direct object manipulation based on explicit instructions, overlooking the nuanced capabilities needed for advanced understanding and response.
To bridge this gap, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [14]. GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes and 97,280 images captured through Kinect Azure and RealSense, featuring 88 objects. Its wide variety of scenes, objects, and grasp poses provides an excellent foundation. In our reasoning grasping dataset, we incorporate reasoning instructions alongside detailed annotations of object parts, and we exclude objects that lack semantic clarity and data associated with invalid grasps. This expanded dataset, now tailored for reasoning grasping tasks, enhances its relevance and applicability for advanced robotic manipulation research.
We provide an overview of our dataset in Fig. 2. In Section V-A, we elaborate on the process of generating reasoning instructions, a novel component of our dataset. Details of object parts annotations are discussed in Section V-B.
@@ V-A Reasoning Instruction Generation
As illustrated in Fig. 3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and meticulous manual review. The generated instructions enable robots to grasp based on nuanced, context-rich queries.
@@ V-A2 Automated Generation with GPT-4
The next step in instruction generation involved utilizing a customized variant of the GPT-4 model. This model was tasked with generating indirect questions and instructions for various objects and their respective parts. The specific prompt provided to the GPT was formulated to elicit responses that allude to the uses or features of an object, without directly naming it. The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks. These are crafted to be specific, relevant, and direct yet subtly phrased, maintaining clarity and functionality.
@@ V-A3 Manual Review and Refinement
Post-generation, each set of GPT-generated instructions underwent a rigorous manual review and refinement process. This step was crucial to ensure the instructions met our standards for subtlety, creativity, and contextual relevance. The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts. This semi-automated approach balanced the efficiency of machine-generated content with the nuanced understanding and judgment of human experts.

### v2

@@ I Introduction
Robotic grasping has long been a subject of extensive study. The utilization of CNN-based neural networks has demonstrated efficiency in generating high-quality grasping poses from visual input [31, 21, 18, 23, 33, 27]. One significant limitation of these methods is the lack of scene understanding. The robot cannot identify the objects they are grasping. Recent methods have introduced language-guided grasping with instructions such as “I want a stapler” [6] or “Pick the food box in front of the ball” [42]. However, those methods will struggle when dealing with implicit instructions, where the desired object is not explicitly named. For example, if a user says “I need to drink water”, an intelligent assistant robot should identify and grasp a cup from a selection of household items, although “cup” does not appear in the user’s instruction. Such scenarios are common in real-world settings and the solutions to them remain open questions.
@@
Recently, the advance of large language models like ChatGPT [32] has introduced new possibilities for language reasoning. Beyond text, numerous studies have leveraged multi-modal Large Language Models (LLMs) for visual understanding [50, 11, 24], enabling these models to interpret and generate responses based on both textual and visual inputs. However, these approaches focus on visual comprehension and text generation, lacking the capability to generate robotic actions. While there have been efforts to integrate LLMs with robotics [2, 16, 20, 36, 43, 17], these concentrate on high-level planning.
In this work, we introduce the novel task of Reasoning Grasping, where a robot determines grasp poses based on users’ “implicit” instructions. “Implicit” instructions are language instructions that do not specify the name of the grasp target. A formal definition of reasoning grasping and implicit instruction is given in Sec III.
@@
To address the reasoning grasping task, we propose a model that directly output grasp poses according to image inputs and language instructions in cluttered environments. Specifically, this model leverages a multi-modal LLM to interpret both images and instructions. We employ a special token strategy to identify tokens that are relevant to the grasping target (i.e., names of target objects). Subsequently, embeddings of these identified tokens are used to generate accurate grasping poses. To train this model, we extend the GraspNet-1 billion dataset [13] with diverse implicit instructions and object part grasping.
@@
  • We introduce a novel task of reasoning grasping, which directs robots to grasp objects based on implicit instructions.
  • We propose a multi-modal LLM for reasoning grasping. Our experiments demonstrate the model’s promising performance.
  • We have developed a unique dataset for the reasoning grasping task. It includes 64 objects, 109 parts, 1730 reasoning instructions, and around 100 million grasping poses. This dataset provides a comprehensive resource for training and evaluating models on reasoning grasping tasks.
@@ II-A Robotic Grasping
Robotic grasping has traditionally relied on analytical methods [28, 10, 35], which focus on understanding the geometry of objects or analyzing contact forces. Convolutional neural networks (CNNs) based methods [31, 21, 18, 23, 33, 27] utilize large datasets of labeled grasping examples [19, 8, 13] to train models to predict grasps based on visual input. However, a critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions. They do not inherently know what they are grasping, which restricts their application in more dynamic, human-centric environments.
Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [26, 38, 40, 5, 46, 37, 45, 47, 49, 1]. This shift allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios. For example, Tziafas et al. focus on referring grasp synthesis, predicting grasp poses for referred objects through natural language in cluttered scenes [42]. Chen et al. proposed to learn from visual and language features jointly and predict the 2D grasp box from the RGB images [6].
@@ II-B LLMs for robotics
Recent breakthroughs in Large Language Models (LLMs) have been impressive [9, 4, 32, 41] in understanding and generating human-like text. Multi-modal LLMs can seamlessly integrate with other modalities, such as vision, enhancing their proficiency in tasks like visual understanding [24]. There are also a lot of efforts integrating LLMs with robotics. Many studies [2, 16, 20] have integrated LLMs into closed-loop planning structures, decomposing language-conditioned long-horizon tasks into small steps. Yet, the gap between language instructions and actions still remains. Furthermore, some studies [36, 43, 17] have employed program-like specifications to prompt LLMs, melding planning and action using a predefined library of action functions.
There is also an increase in research focused on utilizing LLMs specifically for robotic grasping. Tang et al. proposed GraspGPT to use LLMs to generate semantic knowledge and then selected task-oriented grasp pose from grasp candidates [39]. Mirjalili et al. proposed LAN-grasp to identify the optimal part of an object for grasping with LLM, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [29]. Compared to GraspGPT and LAN-grasp, our method is distinguished by its ability to operate in cluttered scenes and its end-to-end training framework, while GraspGPT and LAN-grasp utilize modular frameworks and depend heavily on other pre-trained grasp detection models.
@@ III Reasoning Grasping
In this section, we define the task of Reasoning Grasping, which aims to generate a robotics grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$. In the reasoning grasping task, instructions $t$ are not limited to being straightforward, they can also be “implicit”. We define implicit instructions as language instructions that do not explicitly specify the name of the grasp target but provide relevant contextual cues or indicate users’ intent implying the grasp target. For example, implicit instructions may include cues such as shapes, colors, or other attributes of grasp targets, or they may describe users’ needs or intent implying the grasp targets based on their functionalities.
@@ IV Reasoning Grasping via Multi-modal LLMs
In this section, we present an innovative model for reasoning grasping, which combines conventional robotic grasping in [21] and reasoning capabilities from multi-modal LLMs. Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs into the robotic grasping prediction process.
This model addresses two key challenges: firstly, it addresses the adaptation of pre-trained multi-modal LLMs for efficient training in downstream tasks. Although LLMs have strong reasoning abilities, they often produce verbose and diverse outputs, making it challenging to directly integrate them into training other tasks, especially within an end-to-end framework. Secondly, the model addresses the combination of textual outputs from LLMs and image-based robotic grasping systems. This integration is crucial as it ensures that the rich linguistic context is effectively utilized to refine the subsequent grasping prediction.
@@ IV-A Architecture
The proposed model framework is illustrated in Fig. 1. As for multi-modal LLMs, we utilize the pre-trained LLaVA [24] enhanced with LoRA [15] fine-tuning techniques. LLaVA integrates a CLIP [34] visual encoder with LLaMA [41], an open-source Large Language Model comparable in performance to GPT-3 [4]. And LoRA, known for its computational efficiency, is incorporated into both the projection layer and all linear layers of multi-modal LLMs.
@@
The key role of multi-modal LLMs in the proposed model is to interpret both instruction $t$ and image $v$, then accurately identify the grasping target. Such a target could be an object or, more challenging, a specific part within cluttered scenes. To address the verbosity of LLM outputs, we introduce a special token, i.e., [SPT], strategically placed around the names of the grasping target in the LLM textual outputs. For instance, [SPT] grasping target name [SPT]. The purpose of introducing this special token is to identify and isolate the crucial tokens, specifically, the names of the grasping target, from the textual outputs generated by LLMs.
@@
Once the grasping target is identified using the special token, we proceed to extract the last-layer embeddings of its corresponding tokens. It is important to note that there could be multiple tokens associated with the identified grasping target, and we extract embeddings for all identified tokens. These extracted embeddings encapsulate the essential linguistic information related to the grasping target. We then compute the average of these embeddings denoted as $h_{avg}$, which is passed through a projection network to obtain the feature $f$. This feature $f$ is then fed into the image-based grasping detection process, guiding the prediction of the grasp pose.
@@
The grasping detection in our model, rooted in the CNN-based robotic grasping framework in [21], takes n-channel images as inputs. The input image is first passed through three convolutional layers and five residual layers. The extracted visual features are concatenated with the embeddings obtained from the reasoning module. Following this, we use the convolutional transpose operation to up-sample this concatenated feature, increasing its size to match the input image size. The model subsequently generates four output images, each representing different aspects of the grasp pose $g$: the grasp quality score $q$, the rotation angle expressed in $\cos 2\theta$ and $\sin 2\theta$, and the required width $w$ for the end-effector. The grasp central coordinate $(x,y)$ is determined from the output image of the grasp quality score. This is obtained by selecting the point in the image with the highest quality score. Notably, the depth image $v_{d}$ is an optional input modality for the grasping prediction, depending on data availability.
@@ IV-C Training Strategy
The proposed method leverages the pre-trained LLaVA model, with the LoRA technique for fine-tuning during the training phase. With LoRA fine-tuning, the multi-modal LLM can generate the grasp target within just a few epochs. Based on this observation, our training strategy involves freezing the parameters of the LoRA-enhanced model after the initial epochs. Specifically, from the third epoch, we set the parameter $\lambda_{t}$ in the loss function to 00 and freeze the parameters of the LoRA-enhanced model. Our training then focuses exclusively on other parameters outside the LoRA framework. This approach is designed to reduce the training time and expedite the convergence of the model, particularly in terms of grasping prediction.
@@
Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [13] with additional annotations for part-level grasping and reasoning instructions.
@@ V Reasoning Grasping Dataset
To train the model for reasoning grasping tasks, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [13]. GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes, 88 objects, and 97280 images. Its wide variety of scenes, objects, and grasp poses provides an excellent foundation. We extend the GraspNet-1B dataset with reasoning instructions as well as object parts grasping annotations. We provide some examples in our dataset in Fig. 2.
@@ V-A Reasoning Instruction Generation
As illustrated in Fig. 3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and manual review. The generated instructions enable robots to grasp based on nuanced, context-rich queries.
@@ V-A2 Automated Generation with GPT-4
The next step in instruction generation involved utilizing a customized variant of the GPT-4 model. This model was tasked with generating indirect questions and instructions for various objects and their parts. The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks. These are crafted to be specific, relevant, and direct for grasping tasks.
@@ V-A3 Manual Review and Refinement
Post-generation, each set of GPT-generated instructions underwent a manual review and refinement process. The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts. This semi-automated approach balanced the efficiency of machine-generated content and judgment of human experts.